# Nonlinear dynamics of confined cell migration—modeling and inference

### v1

@@ III.5.1 Perturbing $\beta$, the magnitude of noisy activity
$\beta(t)\sim\mathrm{Gaussian}(\mu_{\beta},\sigma_{\beta})$ is a random variable that sets the magnitude of the stochastic patches we add to cell polarity. We find that our model predicts limit-cycles on two-state and stationary behavior on rectangular micropatterns relatively robustly against variations in both $\mu_{\beta}$ (Fig. 6) and $\sigma_{\beta}$ (Fig. 6). However, there are a few interesting features that can appear within the larger limit-cycle dynamics. For the two cases where $\mu_{\beta}=1.4\mathrm{min}^{-1}$ and $\sigma_{\beta}=2.8\mathrm{min}^{-1}$, a new behavior emerges—in addition to the limit-cycles—that is confined to the interior of the bridge. This suggests that cells that happen to enter the bridge of the two-state micropattern with near-zero velocity may become trapped there, and either perform a small limit-cycle or reach a stationary state. These behaviors at the bridge are ones we are slightly skeptical about observing in experiments akin to [14], in part because our model does not have an explicit representation of the nucleus, while [14] use the cell’s nucleus to parameterize $x$. We would expect the nucleus to not be stably trapped exactly in the middle of the micropattern.
Lastly, we note that for $\sigma_{\beta}=6.6\mathrm{min}^{-1}$, the limit-cycles do not reach from the left basin to the right basin, but instead the streamlines loop around a narrower region. This corresponds to an oscillation of the cell center of mass around the bridge in our plot of $F(x,v)$. When we inspect our movies of cell trajectories, we see that in this case, cells quickly re-enter the bridge once they have entered a basin (see Movie M6).
We varied $\mu_{\beta}$ and $\sigma_{\beta}$ over a range set by monitoring the behavior of the cell. For values lower than $\mu_{\beta}=.63\mathrm{min}^{-1}$ and $\sigma_{\beta}=2.8\mathrm{min}^{-1}$, there isn’t a sufficient buildup of polarity, and we do not observe the cells to hop. Similarly, for higher values, we see many cell body ruptures as polarity becomes large enough to split one cell into many. The perturbations in this section show that our predictions are not overly sensitive to the magnitude of the activity patch so long as $\beta(t)$ is within a reasonable range.
@@ IV Discussion
Within our model, we can switch cells between showing limit-cycle dynamics and bistability either by changing physical parameters like cell size or tension, or by changing polarity dynamics like the size of protrusion fluctuations. Our results indicate that these distinct motility patterns can reflect quantitative changes in the cell, rather than requiring fundamentally different assumptions in how we model MCF10A and MDA-MB-231 cells. Consistent with this idea, it was later found that subpopulations of cancerous cells, derived from their inherent cell-to-cell variability, can also exhibit distinct motility behaviors [53]. Additionally, the expression of the tumor-suppressing micro RNA miR-200c in MDA-MB-231 cells has recently been shown to induce a switch in their motility behavior on the two-state micropattern, with cells expressing miR-200c becoming bistable [54]. While miR-200c has many targets and its impact on cell migration is complex, there is evidence that its expression can lead to increased cell stiffness [55, 56] and decreased cell size [44]. Taken together, these findings complement our predictions that dramatic, qualitative changes in cell behavior can possibly be explained by simple, quantitative changes in the cell.
@@
If the key difference between MCF10A and MDA-MB-231 cells were the cell spread area, which we have shown is sufficient to explain a transition between limit cycle and bistable behaviors, this could be easily tested through both quantification of cell spread areas and variation of the micropattern size. Within our model, simulated cells with similar $A_{\textrm{cell}}/A_{\textrm{basin}}$ have similar dynamics, assuming all other parameters are held constant. To establish further confidence in our findings and probe the limits of our model, we would explore motility in other geometries. For example, there are experimental observations for MDA-MB-231 cells on spiked ratchets [3], as well as various two-state geometries with circular-square and circular-triangular basins [57], all of which provide rich context for further comparisons. Further experimental studies will also be necessary to validate our suggestion of cell properties as predictors of cell motility.
@@
In our application of the data-driven approach of [14], we discovered that uncertainties in the inferred nonlinear dynamics—driven by finite sampling—can be significant. This is despite our large dataset of 4800 simulations spanning for 48 hours, providing roughly 24x as much trajectory time as observed experimentally. This means a single streamline plot should be taken with some skepticism. For instance, the bistable dynamics seen in Fig. 6 with $\sigma=12{\mu\mathrm{m}}$ generates a streamline plot that is slightly asymmetric at the center, and has a small domain of attraction at the center. These features, though, are not robust to sampling (see Appendix E). We suggest that the bootstrap approach we take in Appendix E should be used as an essential robustness check.
@@ Appendix B Constructing acceleration footprints from simulated trajectory data
Appendix C A dynamically-evolving cell boundary is required for bipolar morphology
@@
Appendix D Predicted motility dynamics in response to simplifying model knockouts
Figure S2: Predicted motility dynamics of our default cell on S2 the two-state and S2 the rectangular micropatterns with the filopodial mechanism absent, i.e., $p(\bm{\theta})=\mathbb{P}(\bm{\theta})$.
@@ Appendix D Predicted motility dynamics in response to simplifying model knockouts
Figure S4: Predicted motility dynamics of our default cell on S4 the two-state and S4 the rectangular micropatterns with the perimeter-induced degradation of polarity absent, i.e., $s_{0}\to\infty$.
Appendix E Showcasing the uncertainty of phase-space separatrices
@@
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
@@
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
@@
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
@@
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
Figure S8: Overview of variability in motility for $\mu_{\beta}=1.4\mathrm{min}^{-1}$ perturbation on the two-state geometry presented in Fig. 6.
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
Figure S9: Overview of variability in motility for $\sigma_{\beta}=2.8\mathrm{min}^{-1}$ perturbation on the two-state geometry presented in Fig. 6.
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
@@
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
@@
Refer to caption (a) Refer to caption (b)
Refer to caption (c)
@@
Appendix F Legend for the attached simulation movies
@@
    Movie 6: Default, noisy cell with $\sigma_{\beta}=6.6\mathrm{min}^{-1}$ migrating on the two-state micropattern

### v2

@@ III.5.1 Perturbing $\beta$, the magnitude of noisy activity
$\beta(t)\sim\mathrm{Gaussian}(\mu_{\beta},\sigma_{\beta})$ is a random variable that sets the magnitude of the stochastic patches we add to cell polarity. We find that our model predicts limit-cycles on two-state and stationary behavior on rectangular micropatterns relatively robustly against variations in both $\mu_{\beta}$ (Fig. 6) and $\sigma_{\beta}$ (Fig. 6). However, there are a few interesting features that can appear within the larger limit-cycle dynamics. For the two cases where $\mu_{\beta}=10$ and $\sigma_{\beta}=22.5$, a new behavior emerges—in addition to the limit-cycles—that is confined to the interior of the bridge. This suggests that cells that happen to enter the bridge of the two-state micropattern with near-zero velocity may become trapped there, and either perform a small limit-cycle or reach a stationary state. These behaviors at the bridge are ones we are slightly skeptical about observing in experiments akin to [14], in part because our model does not have an explicit representation of the nucleus, while [14] use the cell’s nucleus to parameterize $x$. We would expect the nucleus to not be stably trapped exactly in the middle of the micropattern.
Lastly, we note that for $\sigma_{\beta}=52.5$, the limit-cycles do not reach from the left basin to the right basin, but instead the streamlines loop around a narrower region. This corresponds to an oscillation of the cell center of mass around the bridge in our plot of $F(x,v)$. When we inspect our movies of cell trajectories, we see that in this case, cells quickly re-enter the bridge once they have entered a basin (see Movie M6).
We varied $\mu_{\beta}$ and $\sigma_{\beta}$ over a range set by monitoring the behavior of the cell. For values lower than $\mu_{\beta}=5$ and $\sigma_{\beta}=22.5$, there isn’t a sufficient buildup of polarity, and we do not observe the cells to hop. Similarly, for higher values, we see many cell body ruptures as polarity becomes large enough to split one cell into many. The perturbations in this section show that our predictions are not overly sensitive to the magnitude of the activity patch so long as $\beta(t)$ is within a reasonable range.
@@ III.5.3 Perturbing $\tau_{f}$, the frequency of noisy activity
Balance is required between some pairs of parameters to generate successful migration We also study the variation of pairs of model parameters simultaneously. We find that there is a balance required between terms that tend to promote cell polarization, and those that degrade it. For instance, if the generation of protrusion pulses is made significantly slower, increasing $\tau_{f}$, cells will not be able to generate enough force to actually cross from one basin to the other, and streamlines will be disconnected (Fig. S1.) However, if the polarity degradation time $\tau$ is also increased, the limit cycle behavior can be recovered. Similarly, decreasing $\tau_{f}$ can be compensated by increasing the amplitude of polarity patches $\mu_{\beta}$ (Fig. S2).
@@ III.6 Can we simplify our model assumptions?
We could also consider other potential model variants. For instance, it would be natural to consider eliminating our micropattern-sensing terms in Eq. 6 and replacing them with, e.g. an energy of adhesion to the micropattern [26, 21, 27]. This would then create an energy barrier for the cell leaving the micropattern. This energy barrier would successfully confine the cell and prevent it from escaping the micropattern, but this variant of the model would not suffice to repolarize the cell – there would be no inhibition of polarity when the cell hit the micropattern edge, preventing it from turning around. Within our model, this would lead to an absence of limit cycles. We believe the role of cell-substrate adhesion to have both an important mechanical effect as well as a signaling (repolarization) effect. Adhesive energy, in combination with alternate models of polarity, might be sufficient to create limit cycles, though this is speculative. Within the context of our current work, we do not treat this adhesive energy, since it is not necessary to recapitulate our core results.
Figure 9: Simulation ($N=2400$) and experiment of cell center of mass distributions in geometries of different shape. Experiment is edited from [62]. As in [62], areas of basins in the asymmetric shapes may differ slightly. Our reported percentages are the fraction of the time the cell spends in one basin, weighted by the area of the basin, $P_{\textrm{left}}=\frac{\left(\textrm{time in left}\right)/A_{\textrm{left}}}{\left(\textrm{time in left}\right)/A_{\textrm{left}}+\left(\textrm{time in right}\right)/A_{\textrm{right}}}$ as in [62].
III.7 Generalization to other micropattern geometries
In a follow-up to [14], Fink et al. studied the effect of varying micropattern shape and size, quantifying the distributions of occupancies of MDA-MB-231 cells in each micropatterned “basin” [62]. We originally tuned our model to reproduce the qualitative features of MDA-MB-231 observed in [14]. How successful is our model at predicting these experiments? We show simulations of cell motility in analogous geometries in Fig. 9. We have not quite matched the experimental geometries – as with our earlier simulations, we find that our cells struggle to move through the absolute narrowest bridges without breaking, so we have chosen bridges to be wide enough for cell to travel through easily. We find in Fig. 9 that our model correctly predicts the side of the micropattern which the cell spends the most time on in eight out of the nine geometries tested. This includes both relatively trivial cases – we find that the cell will spend more time in the larger-area basin as the basin size is increased (Fig. 9a) – but also less intuitive results. For instance, we correctly predict that cells will spend more time in the vertically-oriented rectangular basin, and less in the horizontally-oriented rectangular basin (Fig. 9b). This deviation is smaller than found in the experiment, though it is larger than the presumably statistical asymmetry for the symmetric rectangle-rectangle geometry. We also see that the histograms for the vertical-rectangular vs horizontal-rectangular case are notably asymmetric, as found in the experiment. Cells also spend more time in the triangular basin than a square basin (Fig. 9c) when normalized by area. The only distinct failure is the square-diamond geometry (Fig. 9c), in which we predict more localization to the square, but experiments observe slightly more localization to the diamond.
Overall, we view this as a relatively good performance in generalization. We have a few possible hypotheses for the failure in the square-diamond micropattern. First, this discrepancy with experiment may reflect simply that we have only roughly fit our parameters – we originally sought to qualitatively reproduce the key features of [14], and did not tune most of our parameters beyond that necessary to reproduce the key features. We note in particular that in this square-diamond micropattern, cells can develop three independent fronts (see Fig. 4A of [62]). Our model does not commonly do this – perhaps unsurprisingly, since as we discussed earlier, the coexistence of multiple fronts is an emergent property of our model and one that may require some degree of parameter tuning, based on previous work [50]. We suspect this behavior is not constrained by the requirement that our model reproduces the results of [14]. Another possibility is that the asymmetry in the square-diamond pattern just arises from statistical fluctuations in the experiment. The experimental square-diamond asymmetry is one of the smallest effects observed in [62] (asymmetry of $\sim 2$ times the standard error). Given the multiple comparisons made here and in [62] it is not too surprising that one of these many comparisons would vary by this magnitude.
While [62] did not show streamline plots, we can easily compute $F(x,v)$ and its streamlines (Fig. S7); these predictions could be compared with data generated by [62]. However, we do not see immediate qualitative changes in the streamline plots between the symmetric and asymmetric streamline plots – the changes are primarily quantitative.
@@ IV Discussion
Within our model, we can switch cells between showing limit-cycle dynamics and bistability either by changing physical parameters like cell size or tension, or by changing polarity dynamics like the size of protrusion fluctuations. Our results indicate that these distinct motility patterns can reflect quantitative changes in the cell, rather than requiring fundamentally different assumptions in how we model MCF10A and MDA-MB-231 cells. Consistent with this idea, it was later found that subpopulations of cancerous cells, derived from their inherent cell-to-cell variability, can also exhibit distinct motility behaviors [64]. These findings complement our predictions that dramatic, qualitative changes in cell behavior can possibly be explained by simple, quantitative changes in the cell. Expression of the tumor-suppressing micro RNA miR-200c in MDA-MB-231 cells has recently been shown to induce a switch in their motility behavior on the two-state micropattern, with cells expressing miR-200c becoming bistable [65]. While miR-200c has many targets and its impact on cell migration is complex, it does modify both cell stiffness and size. However, the dynamics here are somewhat complex – its expression can lead to decreased cell stiffness [66] and in different assays miR-200c induction increased cell spread [66] but miR-200 family knockout increased cell size and decreased polarity [54]. If miR-200c induction decreased spread in [65], that could lead to bistability, according to our model. However, perturbations of polarity can also lead to switches between bistability and limit cycles within our model.
To what extent can the parameter variations we have studied be realized in experiment? Past experiments have studied drug perturbations including inhibiting the Rho/ROCK pathway, preventing actin polymerization with latrunculin, and inhibiting myosin contractility with blebbistatin [63], finding that these can disrupt the nonlinear dynamics as visualized in $F(x\to 0,v)$. Within our model, we would expect the application of blebbistatin – decreasing myosin contractility – to decrease our value of $\gamma$, which we see can induce significant perturbations to nonlinear dynamics (Fig. 4. Latrunculin, disrupting actin polymerization, would definitely alter our protrusion strength $\alpha$. We have not seen particularly interesting effects from changing $\alpha$ other than changing the scale of cell speeds, etc. However, F-actin also has roles in cell polarity and can feed back on Rho GTPases [67] – so latrunculin treatment may also alter our polarity parameters like $\tau_{f}$, $\mu_{\beta}$, etc. Another natural approach would be to try eliminating filopodia, as we studied in Fig. S4. This could be done with, e.g. knockdown of myosin X [68] or formins. However, we caution against overinterpretation of these drug assays – off-target effects [69] and feedback loops can confound the apparent results. Another possibility would be to change the concentration of fibronectin on the micropattern. Within our model, this would change the contrast between the micropattern and the region outside – essentially lowering our value of $\chi$. We have performed simulations lowering $\chi$, and find that there are relatively few new nonlinear features until $\chi$ is small enough that the cell is not well-confined to the micropattern. This intervention, too, may not be straightforward to interpret. The effect of substrate adhesiveness on cell speed is non-monotonic [70, 71], and this non-monotonicity depends on actomyosin contractility [70] and may be complicated to model [72]. In addition, as we are modeling primarily the signaling effect of cell-substrate adhesion, we are not as confident in predicting a dose-response between fibronectin concentration and cell polarization. We view changes in micropattern shape as the intervention most likely to yield agreement between simulation and experiment without the need to introduce additional mechanisms.
@@
If the key difference between MCF10A and MDA-MB-231 cells were the cell spread area, which we have shown is sufficient to explain a transition between limit cycle and bistable behaviors, this could be easily tested through both quantification of cell spread areas and variation of the micropattern size. Within our model, simulated cells with similar $A_{\textrm{cell}}/A_{\textrm{basin}}$ have similar dynamics, assuming all other parameters are held constant, and cells will transition to a bistable state as the basin size is increased (Fig. S8).
@@
In our application of the data-driven approach of [14], we discovered that uncertainties in the inferred nonlinear dynamics—driven by finite sampling—can be significant. This is despite our large dataset of 4800 simulations of 48 hours in length, providing roughly 24x as much trajectory time as observed experimentally. This means a single streamline plot should be taken with some skepticism. For instance, the bistable dynamics seen in Fig. 6 with $\sigma=12{\mu\mathrm{m}}$ generates a streamline plot that is slightly asymmetric at the center, and has a small domain of attraction at the center. These features, though, are not robust to sampling (see Appendix G). We suggest that the bootstrap approach we take in Appendix G should be used as an essential robustness check.
How does our model relate to other models of two-state motility, and can it be discriminated from them? Recent work from the Broedersz group and collaborators [63, 73] has hypothesized that cells may be able to polarize while in the constricted region of the micropattern, while being unpolarized in the basins. In [63], this assumption was built into the model, while [73] provided a mechanism based on the alignment of actin filaments. Our model agrees that cells are often polarized while traveling through the constricted region and not as strongly polarized in the basins, though this emerges from our model’s coupling between the cell polarity and the micropattern via sensing the micropattern boundary. In this sense, we expect that our model is in the same broad class as the assumptions of [63], but has a different specific set of assumptions than that of [73]. To distinguish between the two, the most natural test would be to attempt to design a micropattern that created confinement but disrupted actin alignment – possibly by having curved edges on the bridge between basins.
Data and code availability
Code to reproduce this paper has been deposited at Zenodo https://doi.org/10.5281/zenodo.14373111
@@ Appendix B Constructing acceleration footprints from simulated trajectory data
We note that the conditional averaging method we use here, adapted from [14], can have systematic errors due to finite observation time, etc. [61] These can be corrected if the conditional averaging is done using a basis function approach – but the methods do not generalize well to the straightforward binning of [14]. We have chosen to use the binning approach so that our results are as comparable as possible with the original experiments of [14].
Appendix C Varying two parameters simultaneously
Figure S1: Simultaneous perturbation of the timescales $\tau$ and $\tau_{f}$ shows that limit cycle behavior requires that the ratio of these two rates not get too far from our default value. Results are obtained from 2400 simulations.
Figure S2: Simultaneous perturbation of $\tau_{f}$ and $\mu_{\beta}$ shows that a balance between the rate of patch generation and the mean polarity increase at each patch is necessary to gain limit cycle behavior. Results are obtained from 2400 simulations.
Appendix D A dynamically-evolving cell boundary is required for bipolar morphology
@@
Appendix E Predicted motility dynamics in response to simplifying model knockouts
Figure S4: Predicted motility dynamics of our default cell on S4 the two-state and S4 the rectangular micropatterns with the filopodial mechanism absent, i.e., $p(\bm{\theta})=\mathbb{P}(\bm{\theta})$.
@@ Appendix E Predicted motility dynamics in response to simplifying model knockouts
Figure S6: Predicted motility dynamics of our default cell on S6 the two-state and S6 the rectangular micropatterns with the perimeter-induced degradation of polarity absent, i.e., $s_{0}\to\infty$.
Appendix F Streamline plots for different geometries
Here we show streamline plots for alternate geometries, including those showcased in Fig. 9
Figure S7: Predicted streamlines for two alternate two-state geometries are similar to those studied in [62]. Results are obtained from 2400 simulations.
Figure S8: The default cell switches between limit cycle and bistable as the micropattern size is increased.
Appendix G Showcasing the uncertainty of phase-space separatrices
@@
Figure S12: Overview of variability in motility for $\mu_{\beta}=10$ perturbation on the two-state geometry presented in Fig. 6.
Figure S13: Overview of variability in motility for $\sigma_{\beta}=22.5$ perturbation on the two-state geometry presented in Fig. 6.
@@
Appendix H Legend for the attached simulation movies
@@
    Movie 6: Default, noisy cell with $\sigma_{\beta}=52.5$ migrating on the two-state micropattern